You plan to invest $354 per month. Suppose that you are told that the type of investment that you are considering earns an APR of 10%, with monthly compounding. What is the projected value of your investment in 28 years? Assume that you contribute to your investment at the end of each month, beginning one month from today.

Do not round at intermediate steps in your calculation. Round your answer to the nearest dollar. Do not type the $ symbol.
Answer #1

Rate = 10% / 12 = 0.833333%

Number of periods = 28 * 12 = 336

Future value = Annuity * [(1 + r)n - 1] / r

Future value = 354 * [(1 + 0.00833333)336 - 1] / 0.00833333

Future value = 354 * [16.254936 - 1] / 0.00833333

Future value = 354 * 1,830.593088

Future value = 648,030